# On Bayesian Neural Networks with Dependent and Possibly Heavy-Tailed Weights

### v4

@@ 6 Sufficient conditions for (11)
we have $\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(2)}(\mathbb{x}))>0)=1.$ The claim follows by Proposition 6.3. ∎
6.2 Proof of Propositions 6.3, 6.4, 6.5 and 6.6
Proof of Proposition 6.3.
We assume $C_{B}>0$. The proof in the case when $C_{B}=0$ is similar, and therefore omitted. Recall the definition of $\mathbb{K}^{(\ell)}(\mathbb{x})$, $\ell=2,\ldots,L+1$, in the statement of Theorem 4.1, and note that it consists of the sum of three positive semi-definite matrices. Since by assumption (23) we have that $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$ when $a^{(\ell-1)}=0$, the claim follows if we prove that
$\mathbb{E}_{{\rm prior}}[\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))^{\top}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})]$ is positive definite, $\mathbb{P}_{{\rm prior}}$-a.s., (26)
@@
$\mathrm{det}\left(\sum_{j=1}^{\infty}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\right)>0$, $\mathbb{P}_{{\rm prior}}$-a.s. (27)
Proof of (26). By (22), for $(\mathbb{P}_{{\rm prior}})_{\mathbb{K}^{(\ell-1)}(\mathbb{x})}$-almost all $\mathbb{k}$, under the prior, $\zeta_{1}^{(\ell-1)}(\mathbb{x})\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}$ has a Gaussian density (which is obviously strictly positive on $\mathbb{R}^{d}$). So, by Proposition 6.4 we have that the law of $\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\,|\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}$ under $\mathbb{P}_{{\rm prior}}$ and the Lebesgue measure are not singular. By the Lipschitzianity of $\sigma$ and the Gaussianity of $\zeta_{1}^{(\ell-1)}(\mathbb{x})\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}$, all the entries of
@@
are finite. The claim then follows by Proposition 6.5.
Proof of (27). Note that, for $n\geq d$ arbitrarily fixed, by Lemma 3.1 we have
$\Big\{rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))=d\Big\}$ $\equiv\Big\{\sum_{j=1}^{n}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\quad\text{is positive definite}\Big\}$
$\subset\Big\{\sum_{j\geq 1}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\quad\text{is positive definite}\Big\}.$
@@
$\quad\quad\qquad\geq\mathbb{P}_{{\rm prior}}\left(rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))=d\right),\quad\text{for each $n\geq d$.}$ (28)
By Proposition 6.4, for $(\mathbb{P}_{{\rm prior}})_{\mathbb{K}^{(\ell-1)}(\mathbb{x})}$-almost all $\mathbb{k}$, there exist $A=A_{\sigma}\in\mathcal{B}(\mathbb{R}^{d})$ and
@@
$\mathbb{P}_{{\rm prior}}(\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\in C\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=\int_{C}\varphi_{\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}}(y)\mathrm{d}y,\,\,\text{$\forall$ $C\in\mathcal{B}(A)$}$ (29)
@@
$p_{\mathbb{k}}:=\mathbb{P}_{{\rm prior}}(\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\in A\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=\int_{A}\varphi_{\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}}(y)\mathrm{d}y>0.$ (30)
@@
$\mathbb{P}_{{\rm prior}}(D_{n,d}\cap\{rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))\leq d-1\})=0.$ (31)
Thus, by assuming (31),
@@
So (27) follows by (28) and this latter relation if we prove that $\lim_{n\to\infty}\mathbb{P}_{{\rm prior}}(D_{n,d})=1.$ Due to the dominated convergence theorem, in turn, this latter limit holds if
$\lim_{n\to\infty}\mathbb{P}_{{\rm prior}}(D_{n,d}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=1.$
In order to prove this latter relation, we note that since, under $\mathbb{P}_{\rm prior}$, given $\{\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}\}$, the random vectors $\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))$ are independent and identically distributed, recalling the definition of $p_{\mathbb{k}}$ in (30), we have that
@@
Since
$\lim_{n\to\infty}\sum_{k=0}^{d}\binom{n}{k}p_{\mathbb{k}}^{k}(1-p_{\mathbb{k}})^{n-k}=0,$
we have
@@
and the proof is completed.
It remains to prove (31). We start noticing that
@@
{ $\equiv\Biggl\{\text{$\exists$ $j\in\{d,\ldots,n\}$ and $i_{1},\ldots,i_{j}\in\{1,\ldots,n\}$ such that $\sigma(\zeta_{i_{h}}^{(\ell-1)}(\mathbb{x}))\in A$ $\forall$ $h=1,\ldots,j$}$
$\qquad\qquad\qquad\qquad\qquad\text{and $rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x})))))\leq d-1$}\Biggr\}$
$\subseteq\bigcup_{j=d}^{n}\bigcup_{\{i_{1},\ldots,i_{j}\}\subseteq\{1,\ldots,n\}}\Biggl\{\text{ $\sigma(\zeta_{i_{h}}^{(\ell-1)}(\mathbb{x}))\in A$ $\forall$ $h=1,\ldots,j$}$
$\qquad\qquad\qquad\qquad\qquad\text{and $\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))$ are linearly dependent}\Biggr\}.$
Set
$L_{j}:=\{(x_{1},\ldots,x_{j})\in\mathbb{R}^{d\times j}:\,\,x_{1},\ldots,x_{j}\quad\text{are linearly dependent}\}.$
Using the union bound, it follows
@@
$\leq\sum_{j=d}^{n}\sum_{\{i_{1},\ldots,i_{j}\}\subseteq\{1,\ldots,n\}}\mathbb{P}_{{\rm prior}}\Biggl(\text{ $\sigma(\zeta_{i_{h}}^{(\ell-1)}(\mathbb{x}))\in A$ $\forall$ $h=1,\ldots,j$}$
 $\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\text{ and $\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))$ lin. dep.}\Biggr)$
$=\sum_{j=d}^{n}\sum_{\{i_{1},\ldots,i_{j}\}\subseteq\{1,\ldots,n\}}\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in A^{j}\cap L_{j}),$
@@
For each $j=d,\ldots,n$, $\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in A^{j}\cap L_{j}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=0.$ (32)
Define the probability measure
$\eta_{\mathbb{k}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))^{-1}(C))$
$\qquad:=\frac{\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in C\cap A^{j}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})}{\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in A^{j}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})}$
$\qquad\qquad=p_{\mathbb{k}}^{-j}\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in C\cap A^{j}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}),\quad C\in\mathcal{B}(\mathbb{R}^{d\times j}).$ (33)
Here the equality follows by (30) and the fact that, under the prior, given $\mathbb{K}^{(\ell-1)}(\mathbb{x})$, the random vectors $\zeta_{1}^{(\ell-1)}(\mathbb{x}),\ldots,\zeta_{n}^{(\ell-1)}(\mathbb{x})$ are independent and identically distributed. Due to (29), combining this conditional independence with (33) we have that, under $\eta_{\mathbb{k}}$, the joint law of $\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))$ is absolutely continuous with respect to the Lebesgue measure, and so by Proposition 6.6 the random vectors $\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))$ are linearly independent $\eta_{\mathbb{k}}$-a.s.. Therefore
$\eta_{\mathbb{k}}\left(\left(\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))\right)^{-1}(L_{j})\right)=0,$
which gives (32). ∎
Proof of Proposition 6.4.
We will prove that (25) holds with
$A:=\mathbb{R}^{d}\setminus S\quad\text{and}\quad\varphi_{\sigma(X)}(y):=\begin{cases}\sum_{x\in\sigma^{-1}(y)}\frac{\varphi_{X}(x)}{J_{\sigma}(x)}&\text{if $\sigma^{-1}(y)$ is a finite set}\\ +\infty&\text{otherwise}\end{cases}$
@@
and $\varphi_{X}$ is the density of $X$. Let $B\in\mathcal{B}(\mathbb{R}^{d}\setminus S)$ be arbitrarily fixed. By the area formula applied to non-negative and measurable functions $u$ (see Theorem 3.2.2 in [15] and Theorem 2.93 together with the following remark in [2]), we have
$\int_{\mathbb{R}^{d}}u(x)J_{\sigma}(x)\mathrm{d}x=\int_{\mathbb{R}^{d}}\sum_{x\in\sigma^{-1}(y)}u(x)\mathrm{d}y.$ (34)
By taking
$u(x):=\frac{\varphi_{X}(x)}{J_{\sigma}(x)}\mathbb{1}\{\sigma(x)\in B\}$,
we get
@@
$\mathrm{Leb}^{(d)}(T_{0})=0,$ (35)
where $\mathrm{Leb}^{(d)}$ denotes the Lebesgue measure on $\mathbb{R}^{d}$. Therefore, by the absolute continuity of $X$, we have
$\mathbb{P}(X\in\mathbb{R}^{d}\setminus\sigma^{-1}(S))=\mathbb{P}(X\in\mathbb{R}^{d}\setminus T_{1}),$
and we reduce ourselves to prove that $\mathbb{P}(X\in\mathbb{R}^{d}\setminus T_{1})>0$. Since by assumption $X$ has density $\varphi_{X}$ such that $\varphi_{X}>0$ on $\mathbb{R}^{d}$ this latter inequality is equivalent to $\mathrm{Leb}^{(d)}(\mathbb{R}^{d}\setminus T_{1})>0$. In the final part of this proof, we show that this latter claim is guaranteed by the fact that $\sigma$ is non constant and Lipschitz. Reasoning by contradiction suppose that $\mathrm{Leb}^{(d)}(\mathbb{R}^{d}\setminus T_{1})=0$, then for $\mathrm{Leb}^{(d)}$-almost all $x\in\mathbb{R}^{d}$ we have $x\in T_{1}$ and so $J_{\sigma}(x)=0$. Since $\sigma$ acts componentwise on vectors the Jacobian matrix of $\sigma$ at $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is diagonal with diagonal elements $\sigma^{\prime}(x_{i})$, $i=1,\ldots,d$. Therefore, for $\mathrm{Leb}^{(1)}$-almost all $x\in\mathbb{R}$ it holds $\sigma^{\prime}(x)=0$. Since $\sigma$ is Lipschitz continuous this implies that $\sigma$ is constant, which is a contradiction. We finally prove (35). By the area formula (34) with
$u(x):=\frac{1}{J_{\sigma}(x)}\mathbb{1}\{x\in T_{0}\}$
we have
${\rm Leb}^{(d)}(T_{0})$ $=\int_{\mathbb{R}^{d}}\sum_{x\in\sigma^{-1}(y)}\frac{1}{J_{\sigma}(x)}\mathbb{1}\{x\in T_{0}\}\mathrm{d}y=\int_{\sigma(T_{0})}\sum_{x\in\sigma^{-1}(y)}\frac{1}{J_{\sigma}(x)}\mathrm{d}y\leq\int_{S}\sum_{x\in\sigma^{-1}(y)}\frac{1}{J_{\sigma}(x)}\mathrm{d}y=0,$
where the vanishing of the latter integral follows noticing that the Lipschitzianity of $\sigma$ implies ${\rm Leb}^{(d)}(S)=0$, which, in turn, follows by the (co)area formula (34) with the choice $u(x)=\mathbb{1}\{x\in T_{1}\}$ (this is the weak Morse-Sard property, see Remark $(ii)$ after Theorem 1 in [12, Section 3.4.2]). ∎
Proof of Proposition 6.5.

### v5

@@ 6.1. Proof of Theorems 4 and 5
we have $\mathbb{P}_{{\rm prior}}(\mathrm{det}(\mathbb{K}^{(2)}(\mathbb{x}))>0)=1.$ The claim follows by Proposition 1. ∎
6.2. Proof of Propositions 1, 2, 3 and 4
Proof of Proposition 1.
We assume $C_{B}>0$. The proof in the case when $C_{B}=0$ is similar, and therefore omitted. Recall the definition of $\mathbb{K}^{(\ell)}(\mathbb{x})$, $\ell=2,\ldots,L+1$, in the statement of Theorem 1, and note that it consists of the sum of three positive semi-definite matrices. Since by assumption (6.7) we have that $\mathbb{P}_{{\rm prior}}(N_{\ell-1}((0,\infty))=\infty)=1$ when $a^{(\ell-1)}=0$, the claim follows if we prove that
(6.10) $\mathbb{E}_{{\rm prior}}[\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))^{\top}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})]$ is positive definite, $\mathbb{P}_{{\rm prior}}$-a.s.,
@@ 6.2. Proof of Propositions 1, 2, 3 and 4
(6.11) $\mathrm{det}\left(\sum_{j=1}^{\infty}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\right)>0$, $\mathbb{P}_{{\rm prior}}$-a.s.
Proof of (6.10). By (6.6), for $(\mathbb{P}_{{\rm prior}})_{\mathbb{K}^{(\ell-1)}(\mathbb{x})}$-almost all $\mathbb{k}$, under the prior, $\zeta_{1}^{(\ell-1)}(\mathbb{x})\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}$ has a Gaussian density (which is obviously strictly positive on $\mathbb{R}^{d}$). So, by Proposition 2 we have that the law of $\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\,|\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}$ under $\mathbb{P}_{{\rm prior}}$ and the Lebesgue measure are not singular. By the Lipschitzianity of $\sigma$ and the Gaussianity of $\zeta_{1}^{(\ell-1)}(\mathbb{x})\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}$, all the entries of
@@
are finite. The claim then follows by Proposition 3.
Proof of (6.11). Note that, for $n\geq d$ arbitrarily fixed, by Lemma 1 we have
$\Big\{rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))=d\Big\}$
$\qquad\qquad\equiv\Big\{\sum_{j=1}^{n}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\quad\text{is positive definite}\Big\}$
$\qquad\qquad\subset\Big\{\sum_{j\geq 1}T_{j}^{(\ell-1)}\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))\sigma(\zeta_{j}^{(\ell-1)}(\mathbb{x}))^{\top}\quad\text{is positive definite}\Big\}.$
@@
(6.12) $\quad\quad\qquad\geq\mathbb{P}_{{\rm prior}}\left(rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))=d\right),\quad\text{for each $n\geq d$.}$
By Proposition 2, for $(\mathbb{P}_{{\rm prior}})_{\mathbb{K}^{(\ell-1)}(\mathbb{x})}$-almost all $\mathbb{k}$, there exist $A=A_{\sigma}\in\mathcal{B}(\mathbb{R}^{d})$ and
@@
(6.13) $\mathbb{P}_{{\rm prior}}(\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\in C\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=\int_{C}\varphi_{\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}}(y)\mathrm{d}y,\,\,\text{$\forall$ $C\in\mathcal{B}(A)$}$
@@
(6.14) $p_{\mathbb{k}}:=\mathbb{P}_{{\rm prior}}(\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\in A\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=\int_{A}\varphi_{\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x}))\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}}(y)\mathrm{d}y>0.$
@@
(6.15) $\mathbb{P}_{{\rm prior}}(D_{n,d}\cap\{rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))\leq d-1\})=0.$
Thus, by assuming (6.15),
@@
So (6.11) follows by (6.12) and this latter relation if we prove that $\lim_{n\to\infty}\mathbb{P}_{{\rm prior}}(D_{n,d})=1.$ Due to the dominated convergence theorem this latter limit holds if $\lim_{n\to\infty}\mathbb{P}_{{\rm prior}}(D_{n,d}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=1.$ In order to prove this latter relation, we note that since, under $\mathbb{P}_{\rm prior}$, given $\{\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k}\}$, the r.v.’s $\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))$ are i.i.d., recalling the definition of $p_{\mathbb{k}}$ in (6.14), we have that
@@
Since $\lim_{n\to\infty}\sum_{k=0}^{d}\binom{n}{k}p_{\mathbb{k}}^{k}(1-p_{\mathbb{k}})^{n-k}=0,$ we have
@@
and the proof is completed. It remains to prove (6.15). We start noticing that
@@
{ $\equiv\Biggl\{\text{$\exists$ $j\in\{d,\ldots,n\}$ and $i_{1},\ldots,i_{j}\in[n]$ such that $\sigma(\zeta_{i_{h}}^{(\ell-1)}(\mathbb{x}))\in A$ $\forall$ $h\in[j]$}$
$\qquad\qquad\qquad\qquad\qquad\text{and $rk((\sigma(\zeta_{1}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{n}^{(\ell-1)}(\mathbb{x}))))\leq d-1$}\Biggr\}$
$\subseteq\bigcup_{j=d}^{n}\bigcup_{\{i_{1},\ldots,i_{j}\}\subseteq[n]}\Biggl\{\text{ $\sigma(\zeta_{i_{h}}^{(\ell-1)}(\mathbb{x}))\in A$ $\forall$ $h\in[j]$}$
$\qquad\qquad\qquad\qquad\qquad\text{and $rk((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))))\leq d-1$}\Biggr\}.$
Set $L_{j}:=\{(x_{1},\ldots,x_{j})\in\mathbb{R}^{d\times j}:\,\,rk((x_{1},\ldots,x_{j}))\leq d-1\}.$ Using the union bound, it follows
@@
$\leq\sum_{j=d}^{n}\sum_{\{i_{1},\ldots,i_{j}\}\subseteq[n]}\mathbb{P}_{{\rm prior}}\Biggl(\text{ $\sigma(\zeta_{i_{h}}^{(\ell-1)}(\mathbb{x}))\in A$ $\forall$ $h\in[j]$}$
 $\qquad\qquad\qquad\qquad\text{ and $rk(\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\leq d-1$ }\Biggr)$
$=\sum_{j=d}^{n}\sum_{\{i_{1},\ldots,i_{j}\}\subseteq[n]}\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in A^{j}\cap L_{j}),$
@@
(6.16) For each $j=d,\ldots,n$, $\mathbb{P}_{{\rm prior}}((\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x})))\in A^{j}\cap L_{j}\,|\,\mathbb{K}^{(\ell-1)}(\mathbb{x})=\mathbb{k})=0.$
To this aim, recall that, under the prior, given $\mathbb{K}^{(\ell-1)}(\mathbb{x})$, the r.v.’s $\zeta_{1}^{(\ell-1)}(\mathbb{x}),\ldots,\zeta_{n}^{(\ell-1)}(\mathbb{x})$ are i.i.d and so, using (6.13), we have that, under the prior, given $\mathbb{K}^{(\ell-1)}(\mathbb{x})$, the joint law of $\sigma(\zeta_{i_{1}}^{(\ell-1)}(\mathbb{x})),\ldots,\sigma(\zeta_{i_{j}}^{(\ell-1)}(\mathbb{x}))$ is absolutely continuous with respect to the Lebesgue measure. Therefore, if $\mathbb{A}_{j}\in\mathbb{R}^{j\times d}$ is the matrix having these vectors as columns, then the rows of $\mathbb{A}_{j}$ satisfy the hypotheses of Proposition 4 with $r=d$ and $q=j$. Hence $rk(\mathbb{A}_{j})=d$, and we conclude that (6.16) holds. ∎
Proof of Proposition 2.
We will prove that (6.9) holds with
$A:=\mathbb{R}^{d}\setminus S\quad\text{and}\quad\varphi_{\sigma(X)}(y):=\begin{cases}\sum_{x\in\sigma^{-1}(y)}\frac{\varphi_{X}(x)}{J_{\sigma}(x)}&\text{if $\sigma^{-1}(y)$ is a countable set}\\ +\infty&\text{otherwise}\end{cases}$
@@
and $\varphi_{X}$ is the density of $X$. Let $B\in\mathcal{B}(\mathbb{R}^{d}\setminus S)$ be arbitrarily fixed. By the area formula applied to non-negative and measurable functions $u$ (see Theorem 3.2.2 in [15] and Theorem 2.71 in [2], together with Eq. (2.47) thereafter) we have
(6.17) $\int_{\mathbb{R}^{d}}u(x)J_{\sigma}(x)\mathrm{d}x=\int_{\mathbb{R}^{d}}\sum_{x\in\sigma^{-1}(y)}u(x)\mathrm{d}y.$
By taking $u(x):=\frac{\varphi_{X}(x)}{J_{\sigma}(x)}\mathbb{1}\{\sigma(x)\in B\}$, we get
@@
(6.18) $\mathrm{Leb}^{(d)}(T_{0})=0,$
where $\mathrm{Leb}^{(d)}$ denotes the Lebesgue measure on $\mathbb{R}^{d}$. Therefore, by the absolute continuity of $X$, we have $\mathbb{P}(X\in\mathbb{R}^{d}\setminus\sigma^{-1}(S))=\mathbb{P}(X\in\mathbb{R}^{d}\setminus T_{1}),$ and we reduce ourselves to prove that $\mathbb{P}(X\in\mathbb{R}^{d}\setminus T_{1})>0$. Since by assumption $X$ has density $\varphi_{X}$ such that $\varphi_{X}>0$ on $\mathbb{R}^{d}$ this latter inequality is equivalent to $\mathrm{Leb}^{(d)}(\mathbb{R}^{d}\setminus T_{1})>0$. In the final part of this proof, we show that this latter claim is guaranteed by the fact that $\sigma$ is non constant and Lipschitz. Reasoning by contradiction suppose that $\mathrm{Leb}^{(d)}(\mathbb{R}^{d}\setminus T_{1})=0$, then for $\mathrm{Leb}^{(d)}$-almost all $x\in\mathbb{R}^{d}$ we have $x\in T_{1}$ and so $J_{\sigma}(x)=0$. Since $\sigma$ acts componentwise on vectors, the Jacobian matrix of $\sigma$ at $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is diagonal with diagonal elements $\sigma^{\prime}(x_{i})$, $i\in[d]$. Therefore, for $\mathrm{Leb}^{(1)}$-almost all $x\in\mathbb{R}$ it holds $\sigma^{\prime}(x)=0$. Since $\sigma$ is Lipschitz continuous this implies that $\sigma$ is constant, which is a contradiction. We finally prove (6.18). By the area formula (6.17) with $u(x):=\frac{1}{J_{\sigma}(x)}\mathbb{1}\{x\in T_{0}\}$ we have
${\rm Leb}^{(d)}(T_{0})$ $=\int_{\mathbb{R}^{d}}\sum_{x\in\sigma^{-1}(y)}\frac{1}{J_{\sigma}(x)}\mathbb{1}\{x\in T_{0}\}\mathrm{d}y$
$=\int_{\sigma(T_{0})}\sum_{x\in\sigma^{-1}(y)}\frac{1}{J_{\sigma}(x)}\mathrm{d}y\leq\int_{S}\sum_{x\in\sigma^{-1}(y)}\frac{1}{J_{\sigma}(x)}\mathrm{d}y=0,$
where the vanishing of the latter integral follows noticing that the Lipschitzianity of $\sigma$ implies ${\rm Leb}^{(d)}(S)=0$, which, in turn, follows by the (co)area formula (6.17) with the choice $u(x)=\mathbb{1}\{x\in T_{1}\}$ (this is the weak Morse-Sard property, see Remark $(ii)$ after Theorem 1 in [12, Section 3.4.2]). ∎
Proof of Proposition 3.